# DiffLoRA: Generating Personalized Low-Rank Adaptation Weights with Diffusion

### v1

@@ Conclusion
In this work, we have presented DiffLoRA, a novel method for human-centric personalized image generation. Our approach employs a latent diffusion-based hypernetwork framework to predict LoRA weights for specific identity adaptation in the SDXL, enabling tuning-free generation of high-fidelity portraits without extra inference cost. Experimental results show that DiffLoRA outperforms existing methods in text-image consistency, identity fidelity, generation quality, and inference efficiency. Overall, we believe DiffLoRA is the first to utilize diffusion models for LoRA weights generation, paving the way for more adaptive frameworks in diverse architectures.

### v2

@@ Conclusion
In this work, we have presented DiffLoRA, a novel method for human-centric personalized image generation. Our approach employs a latent diffusion-based hypernetwork framework to predict LoRA weights for specific identity adaptation in the SDXL, enabling tuning-free generation of high-fidelity portraits without extra inference cost. Experimental results show that DiffLoRA outperforms existing methods in text-image consistency, identity fidelity, generation quality, and inference efficiency. Overall, we believe DiffLoRA is the first to utilize diffusion models for identity-oriented LoRA weights generation, paving the way for more adaptive frameworks in diverse architectures.